The lifetime T in hours of a system (with some redundancy) has pdf

f_{T}(t) = Cte^{-\frac{t}{2}}1_{[0,\infty )}(t)

What is the probability that the system will last for at least 5 hours?
Answer #1

here for this to be valid:

\int_{0}^{\infty } f(t) dt should be equal to 1

\int_{0}^{\infty } f(t) dt =\int_{0}^{\infty }Cte-t/2 dt =C*(-2te-t/2-4e-t/2) |\infty0 =C*4 =1

C=1/4

therefore P( system will last for at least 5 hour )=P(X>5)=1-P(X<5)

=1-\int_{0}^{5 } f(t) dt =1-\int_{0}^{5 }(1/4)te-t/2 dt =1-(1/4)*(-2te-t/2-4e-t/2) |50 =1-0.7127=0.2873